OpenStudy (anonymous):

write an equation of the line in point-slope form that passes through the two given points. (-2, 15), (9, -18) THANKS!

OpenStudy (anonymous):

y - 2 = -3 (x - 15) y - 2 = -1/3 (x =+ 15) y - 15 = -3 (x + 2) y - 15 = -1/3 (x + 2) @ganeshie8 @dumbcow @Lost??? @adrynicoleb PLEASE HELP